OpenStudy (poetryprincess):

what is the value of 5^3i^9? A.–125i B.–15i C.15i D.125i

OpenStudy (jhannybean):

\[\huge 5^{3\cdot i^9}\]\[\large i^9 = i^{2+2+2+2}+i^1\implies \color{red}{(i^2)(i^2)(i^2)(i^2)}+i^1\]\[\large \color{red}{i^2 =-1}\]\[\large 5^3 = 125\]You can go on from here.

OpenStudy (poetryprincess):

Thanks!

OpenStudy (jhannybean):

No problem :)